Find the value of C such that the graph of 3x + C = 5y has an x-intercept of (−4, 0)
Answer by stanbon(75887)

Find the value of C such that the graph of 3x + C = 5y has an x-intercept of (−4, 0)
------------
Substitute for x and y and solve for "c":
3(-4) + C = 5*0
-12 + C = 0
C = 12
==========
Cheers,
Stan H.
==========